# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ 1.4 The asymptotic Hecke algebra
Lusztig’s a-function $a\colon\tilde{W}\to\mathbb{Z}_{\geq 0}$ is defined such that $a(w)$ is the minimal value such that $\mathbf{q}^{\frac{a(w)}{2}}h_{x,y,w}\in\mathcal{A}^{+}=\operatorname{Spec}\mathbb{C}[\mathbf{q}^{1/2}]$ for all $x,y\in\tilde{W}$.
The $a$-function is constant of two-sided cells of $\tilde{W}$. Obviously, $a(1)=1$, and the $a$-function obtains its maximum, equal to the number of positive roots, on the two-sided cell containing the longest word $w_{0}\in W$. In general, under the bijection between two-sided cells $\mathbf{c}$ of $\tilde{W}$ and unipotent conjugacy classes $u=u(\mathbf{c})$ in $G^{\vee}$ of [affineIV], we have
@@
where $\mathcal{B}_{u}^{\vee}$ is the Springer fibre of $u$. We have $a(w)\leq\ell(w)$ for all $w\in\tilde{W}$.
In [affineII], Lusztig defined an associative algebra $J$ over $\mathbb{C}$ equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ which becomes an isomorphism after taking a certain completion, to be recalled in Section 3.3.1, of both sides. As an abelian group, $J$ is free with basis $\{t_{w}\}_{w\in\tilde{W}}$. The structure constants of $J$ are obtained from those of $\mathbf{H}$ written in the $\{C_{w}\}_{w\in\tilde{W}}$-basis under the following procedure: first, the integer $\gamma_{x,y,z}$ is defined by the condition
@@
where $\mathcal{D}$ is the (finite) set of distinguished involutions [affineII]. The elements $t_{d}$ are orthogonal idempotents in $J$, which decomposes as a direct sum indexed by the two-sided cells of $\tilde{W}$ in the sense of [affineI]. Each left cell, again in the sense of op. cit., contains a single distinguished involution which is the unit in the ring $t_{d}Jt_{d}$. If $\mathbf{c}$ is a two-sided cell, then $J_{\mathbf{c}}$ is a is a unital ring with unit
@@ 1.4.1 Deformations of the group ring
Upon setting $\mathbf{q}=1$, $\mathbf{H}|_{\mathbf{q}=1}$ is isomorphic to $\mathbb{C}[\tilde{W}]$, and so $\mathbf{H}$ is a deformation of the group algebra of $\tilde{W}$.
@@
Let $W=\left\langle 1,s\,\middle|\,s^{2}=1\right\rangle$ be the Weyl group of type $A_{1}$. The Kazhdan-Lusztig $C_{w}$-basis elements are $C_{1}=T_{1}$ and $C_{s}=\mathbf{q}^{-\frac{1}{2}}T_{s}-\mathbf{q}^{\frac{1}{2}}T_{1}$. There are two two-sided cells in $W$, and one can easily check that
@@ 1.5 Representation theory of $\mathbf{H}$ and $J$
We now recall an algebraic version of the Langlands classification. As we do not have access to the notion of absolute value of $\mathbf{q}$, the classical definitions of tempered and discrete-series representations of the corresponding $p$-adic group $G$ are not available to us. However, Kazhdan-Lusztig provide the following algebraic generalization. Let $\kappa=\mathbb{C}(\mathbf{q}^{-\frac{1}{2}})$ and $\bar{\kappa}$ be the algebraic closure of $\kappa$. We write $\mathbf{H}_{\bar{\kappa}}$ for $\mathbf{H}\otimes_{\mathcal{A}}\bar{\kappa}$ and recall another definition of Lusztig’s from [affineIV].
@@
Let $M$ be a $H_{\bar{\kappa}}$-module finite-dimensional over $\bar{\kappa}$. Say that $m\in M$ is an eigenvector if $\theta_{x}\cdot m=\chi_{m}(x)m$ for all dominant $x$ in $X_{*}$. As $\chi_{m}$ is a character of the coweight lattice, it corresponds to an element $\sigma_{m}\in\mathbf{A}^{\vee}(\bar{\kappa})$ in the sense that, for all cocharacters $x$ of $A$, we have
@@
such that for all eigenvectors $m$ of $M$, the element $\sigma_{m}$ is $\mathbf{G}^{\vee}(\bar{\kappa})$-conjugate to
@@
where by abuse of notation we have written $\phi^{\prime}$ again for the base-change to $\bar{\kappa}$.
@@
Next is a generalization of Casselman’s criterion, which as such, requires a choice of dominant weights. The definitions in both [KLDeligneLanglands] and [affineIV] are both given under the choice of positive roots to be those associated to the opposite Borel (the so-called geometric choice of dominance). On the other hand, for purposes of harmonic analysis, Casselman’s criterion as usually stated uses the standard choice of positive roots.
Following [affineIV], we choose a morphism of groups $V\colon\bar{\kappa}^{\times}\to\mathbb{R}$ such that $V(\mathbf{q}^{-\frac{1}{2}})=1$ and $V(a\mathbf{q}^{-\frac{1}{2}}+b)=0$ for all $a\in\mathbb{C}$, $b\in\mathbb{C}^{\times}$.
@@
Let $M$ be any finite-dimensional $\mathbf{H}_{\bar{\kappa}}$-module and fix the geometric notion of dominance for $X_{*}(\mathbf{A})$. We say that $M$ is $V$-geometrically anti-tempered if all eigenvalues $\nu$ of $\theta_{\lambda}$ for all geometrically dominant $\lambda\in X_{*}(\mathbf{A})$ satisfy $V(\nu)\geq 0$. We say that $M$ is $V$-geometrically anti-square-integrable if $V(\nu)>0$.
When $\mathbf{q}$ is specialized a prime power $q\in\mathbb{C}^{\times}$ and $V(z)=\log|z|$, we recover that our anti-tempered for the geometric choice of dominance modules are tempered in the traditional analytic sense after twisting by the Goldman involution, and likewise for square-integrability, by the Casselman criterion.
@@
Let $E$ be a $J$-module. Then $E\otimes_{\mathbb{C}}\kappa$ is a $J\otimes_{\mathbb{C}}\kappa$-module. Hence $\mathbf{H}_{\kappa}$ acts on $E$ via $\phi$. Denote the resulting $\mathbf{H}_{\kappa}$ module by ${}^{\phi}E$.
@@ 1.5.1 Involutions on $\mathbf{H}$
The reason for our choice of conventions, which match those of [affineIII] and [affineIV], is the presence of the Goldman involution and its exchange of temperedness and anti-temperedness in the relationship between $H$-modules and $J$-modules; see Theorem 4 and Lemma 1 below.
@@
The Goldman involution $h\mapsto h^{\dagger}$ is the $\mathcal{A}$-algebra involution of $\mathbf{H}$ defined by setting
$T_{w}^{\dagger}=(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1}.$
Given an $\mathbf{H}$ (or $\mathbf{H}_{\kappa}$)-module $M$, define ${}^{\dagger}M$ to be the same vector space with the $\mathbf{H}$-action twisted by this involution.
Let $\Delta$ be the usual positive system determined by $\mathbf{B}$ and let $\Delta^{-}$ be the geometric system. Then we have
@@
  1. 1.
    $\overline{h^{\dagger}}=(\bar{h})^{\dagger}=j(h)$
    as $\mathcal{A}$-antilinear involutions. In particular, $C_{w}^{\dagger}=j(C_{w})=(-1)^{\ell(w)}C^{\prime}_{w}$.
  2. 2.
    Let $\lambda$ be dominant for $\Delta$ and let $\theta^{\Delta}_{\lambda}$ and $\theta^{\Delta^{-}}_{\lambda}$ denote Bernstein elements with respect to $\Delta$ and $\Delta^{-}$, respectively. Then
    ${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda}.$
@@
We compute
$\overline{\sum_{x}b_{x}T_{x}^{\dagger}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\overline{\mathbf{q}^{\ell(x)}\bar{T}_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}=j\left(\sum_{x}b_{x}T_{x}\right)$
@@
$\left(\overline{\sum_{x}b_{x}T_{x}}\right)^{\dagger}=\left(\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}^{\dagger}\right)^{\dagger}.$
Thus we have $\bar{C_{w}^{\dagger}}=(-1)^{\ell(w)}C^{\prime}_{w}$, whence the first claim.
Next,
${\theta^{\Delta}_{\lambda}}^{\dagger}=(-1)^{\ell(\lambda)}\mathbf{q}^{-\frac{\ell(\lambda)}{2}}\mathbf{q}^{\ell(\lambda)}T_{-\lambda}^{-1}=(-1)^{\ell(\lambda)}\theta^{\Delta^{-}}_{\lambda},$
noting that $-\lambda$ is dominant with respect to $\Delta^{-}$ and so can appear as a label with respect to $\Delta^{-}$. ∎
@@
Fix the geometric notion of dominance for $X_{*}(\mathbf{A})$. Then there are bijections of sets
@@
Moreover, for a simple $J$-module $E$,
@@
    There is a unique two-sided cell $\mathbf{c}=\mathbf{c}(E)$ of $\tilde{W}$ such that $\mathrm{trace}\left({E}\,,{t_{w}}\right)\neq 0$ implies $w\in\mathbf{c}$.
@@
In particular, $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is independent of $\mathbf{q}$, and upon specializing $\mathbf{q}=q$ a prime power, will be a regular function in the twisting character in the setting of the Payley-Weiner theorem for the Iwahori-Hecke algebra of the $p$-adic group $G$, as we will explain in greater detail below.
That is, in our conventions, simple tempered $\mathbf{H}|_{\mathbf{q}=q}$-modules in the sense of harmonic analysis will extend to simple $J$-modules via ϕ()italic-$\phi\circ()^{\dagger}$.
We will comment even further on the necessity of twisting by the Goldman involution in Section 2.4.
1.6 Acknowledgements
The author thanks Alexander Braverman for introducing him to this material and the problem, for suggesting the use of the Plancherel formula, and for helpful conversations. The author also thanks Roman Bezrukavnikov, Itai Bar-Natan, Gurbir Dhillon, Malors Espinosa-Lara, Julia Gordon, George Lusztig, and Matthew Sunohara for helpful conversations. The author thanks Maarten Solleveld for explaining the generalization in [Solleveld] of the results of [fomFormalDegUnip], suggesting the reference [APM], and for careful reading of an early version of this manuscript.

### v6

@@ 1.4 The asymptotic Hecke algebra
Lusztig’s a-function $a\colon\tilde{W}\to\mathbb{Z}_{\geq 0}$ is defined such that $a(w)$ is the minimal value such that $\mathbf{q}^{\frac{a(w)}{2}}h_{x,y,w}\in\mathcal{A}^{+}=\mathbb{C}[\mathbf{q}^{1/2}]$ for all $x,y\in\tilde{W}$.
The $a$-function is constant of two-sided cells of $W_{\mathrm{aff}}$. Obviously, $a(1)=0$, and the $a$-function obtains its maximum, equal to the number of positive roots, on the two-sided cell containing the longest word $w_{0}\in W$. In general, under the bijection between two-sided cells $\mathbf{c}$ of $W_{\mathrm{aff}}$ and unipotent conjugacy classes $u=u(\mathbf{c})$ in $G^{\vee}$ of [affineIV], we have
@@
where $\mathcal{B}_{u}^{\vee}$ is the Springer fibre of $u$. We have $a(w)\leq\ell(w)$ for all $w\in W_{\mathrm{aff}}$.
In [affineII], Lusztig defined an associative algebra $J$ over $\mathbb{C}$ equipped with an injection $\phi\colon\mathbf{H}\mathrel{\hookrightarrow}J\otimes_{\mathbb{C}}\mathcal{A}$ which becomes an isomorphism after taking a certain completion, to be recalled in Section 3.3.1, of both sides. As a $\mathbb{C}$-vector space, $J$ has basis $\{t_{w}\}_{w\in W_{\mathrm{aff}}}$. The structure constants of $J$ are obtained from those of $\mathbf{H}$ written in the $\{C_{w}\}_{w\in W_{\mathrm{aff}}}$-basis under the following procedure: first, the integer $\gamma_{x,y,z}$ is defined by the condition
@@
where $\mathcal{D}$ is the (finite) set of distinguished involutions [affineII]. The elements $t_{d}$ are orthogonal idempotents in $J$, which decomposes as a direct sum indexed by the two-sided cells of $W_{\mathrm{aff}}$ in the sense of [affineI]. Each left cell, again in the sense of op. cit., contains a single distinguished involution which is the unit in the ring $t_{d}Jt_{d}$. If $\mathbf{c}$ is a two-sided cell, then $J_{\mathbf{c}}$ is a is a unital ring with unit
@@
Lemma 1.
Let
$\varphi\colon(W_{\mathrm{aff}},S)\to(W_{\mathrm{aff}},S)$
be a Coxeter group automorphism. Then
  1. 1.
    The map $t_{w}\mapsto t_{\varphi(w)}$ defines a based ring automorphism of $J$, which we also denote $\varphi$.
  2. 2.
    The map $\phi$ is $\varphi$-linear, in the sense that it commutes with the automorphism from the first statement and the automorphism (3).
Proof.
The first statement is [LX, 2.2(g)].
For the second statement, note that $\varphi$ acts on $\mathcal{D}$ and that
$\phi(C_{\varphi(w)})=\sum_{d_{1}\sim_{L}z_{1}}h_{\varphi(w),d_{1},z_{1}}t_{z_{1}}$
and
$\varphi\left(\phi(C_{w})\right)=\sum_{d_{2}\sim_{L}z_{2}}h_{w,d_{2},z_{2}}t_{\varphi(z_{2})}.$
Now if $z_{1}=\varphi(z_{2})$, then $d_{1}=\varphi(d_{2})$, and
$h_{\varphi(w),d_{1},z_{1}}=h_{\varphi(w),\varphi(d_{2}),\varphi(d_{2})}=h_{w,d_{2},z_{2}}$
by (3). ∎
@@ 1.4.1 Deformations of the group ring
Upon setting $\mathbf{q}=1$, $\mathbf{H}|_{\mathbf{q}=1}$ is isomorphic to $\mathbb{C}[\tilde{W}$], and so $\mathbf{H}$ is a deformation of the group algebra of $W_{\mathrm{aff}}$.
@@
Let $W=\left\langle 1,s\,\middle|\,s^{2}=1\right\rangle$ be the Weyl group of type $A_{1}$. The Kazhdan-Lusztig $C_{w}$-basis elements are $C_{1}=T_{1}$ and $C_{s}=\mathbf{q}^{-\frac{1}{2}}T_{s}-\mathbf{q}^{\frac{1}{2}}T_{1}$, and $\mathcal{D}=W$ in this case. There are two two-sided cells in $W$, and one can easily check that
@@ 1.5 Representation theory of $\mathbf{H}$ and $J$
We now recall an algebraic version of the Langlands classification. As we do not have access to the notion of absolute value of $\mathbf{q}$, the classical definitions of tempered and discrete-series representations of the corresponding $p$-adic group $G$ are not available to us. However, Kazhdan-Lusztig provide the following algebraic generalization. Let $\mathcal{K}=\mathbb{C}(\mathbf{q}^{-\frac{1}{2}})$ and $\bar{\mathcal{K}}$ be the algebraic closure of $\mathcal{K}$. We write $\mathbf{H}_{\bar{\mathcal{K}}}$ for $\mathbf{H}\otimes_{\mathcal{A}}\bar{\mathcal{K}}$ and recall another definition of Lusztig’s from [affineIV].
@@
Let $M$ be a $H_{\bar{\mathcal{K}}}$-module finite-dimensional over $\bar{\mathcal{K}}$. Say that $m\in M$ is an eigenvector if $\theta_{x}\cdot m=\chi_{m}(x)m$ for all dominant $x$ in $X_{*}$. As $\chi_{m}$ is a character of the coweight lattice, it corresponds to an element $\sigma_{m}\in\mathbf{A}^{\vee}(\bar{\mathcal{K}})$ in the sense that, for all cocharacters $x$ of $A$, we have
@@
such that for all eigenvectors $m$ of $M$, the element $\sigma_{m}$ is $\mathbf{G}^{\vee}(\bar{\mathcal{K}})$-conjugate to
@@
where by abuse of notation we have written $\phi^{\prime}$ again for the base-change to $\bar{\mathcal{K}}$.
@@
Next is a generalization of Casselman’s criterion, which as such, requires a choice of dominant weights. Following [KLDeligneLanglands] (see the proof of Prop. 1.6 of loc. cit.) and [affineIV, Section 1.6], we choose the positive roots to be those occurring in $\mathfrak{g}/\mathfrak{b}$.
Following [affineIV], we choose a morphism of groups $V\colon\bar{\mathcal{K}}^{\times}\to\mathbb{R}$ such that $V(\mathbf{q}^{\frac{1}{2}})=1$ and $V(a\mathbf{q}^{\frac{1}{2}}+b)=0$ for all $a\in\mathbb{C}$, $b\in\mathbb{C}^{\times}$.
@@
Let $M$ be any finite-dimensional $\mathbf{H}_{\bar{\mathcal{K}}}$-module. We say that $M$ is $V$-tempered if all eigenvalues $\nu$ of $\theta_{\lambda}$ for all dominant $\lambda\in X_{*}(\mathbf{A})$ satisfy $V(\nu)\leq 0$.
@@
Let $E$ be a $J$-module. Then $E\otimes_{\mathbb{C}}\mathcal{K}$ is a $J\otimes_{\mathbb{C}}\mathcal{K}$-module. Hence $\mathbf{H}_{\mathcal{K}}$ acts on $E$ via $\phi$. Denote the resulting $\mathbf{H}_{\mathcal{K}}$ module by $^{\phi}E$.
@@ 1.5.1 Involutions on $\mathbf{H}$
For $x\in W_{\mathrm{aff}}$, let $\omega(x)\in\pi_{1}(G)$ label the $W\ltimes\mathbb{Z}\Phi^{\vee}$-coset of $W_{\mathrm{aff}}$ containing $x$, and write $\omega(x)=\omega(x)_{f}\omega(x)_{t}\in W\ltimes X_{*}$. Then every $y\leq x$ is also in the coset of $\omega(x)$.
@@
The reason for our choice of conventions, which differ slightly from those of [affineIII] and [affineIV], is the presence of the involution $\dagger$ and its exchange of temperedness and anti-temperedness in the relationship between $H$-modules and $J$-modules; see Theorem 4 and Lemma 2 below.
@@
  1. (a)
    Define the $\mathcal{A}$-algebra involution $^{\dagger}(-)$ of $\mathbf{H}$ by setting
    $^{\dagger}T_{w}=(-1)^{\ell(w_{f})}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1},~~~w=w_{f}\lambda\in W_{\mathrm{aff}}.$
    Note that the sign factor depends only on $w_{f}$.
  2. (b)
    Let $h\mapsto{}^{*}h$ be the $\mathcal{A}$-involution defined in terms of the Bernstein presentation of $\mathbf{H}$: by
    $^{*}T_{w}=(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1},~~~w\in W$
    and
    $^{*}\theta_{\lambda}=\theta_{\lambda}^{-1},~~~\lambda\in X_{*}.$
  3. (c)
    ([PrasadDuke, Section 5.a].) Let $\bar{\kappa}$ be the $\mathcal{A}$-linear involution defined by
    $\bar{\kappa}(T_{w})=T_{\bar{\kappa}(w)}$
    induced by the Coxeter group automorphism given by $\bar{\kappa}(s)=w_{0}sw_{0}$ for $s\in S\setminus\left\{s_{0}\right\}$ and $\bar{\kappa}(s_{0})=s_{0}$, so that $\bar{\kappa}(\lambda)=-w_{0}(\lambda)$ for $\lambda\in X_{*}$; equivalently
    $\bar{\kappa}(T_{w})=T_{w_{0}ww_{0}^{-1}},~~~w\in W,~~~\bar{\kappa}(\theta_{\lambda})=\theta_{-w_{0}(\lambda)},~~~\lambda\in X_{*}.$
  4. (d)
    ([Barbasch-Moy, Section 5].) Let $h\mapsto{}^{\bullet}h$ be the $\mathcal{A}$-linear anti-involution defined by
    $^{\bullet}T_{w}=T_{w^{-1}},~~~w\in W,~~~{}^{\bullet}\theta_{\lambda}=\theta_{\lambda},~~\lambda\in X_{*}.$
  5. (e)
    ([Barbasch-Moy, Section 5].) Let $h\mapsto{}^{\star}h$ be the $\mathcal{A}$-linear anti-involution defined by
    $^{\star}T_{w}=T_{w^{-1}},~~~w\in W,~~~{}^{\star}\theta_{\lambda}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}T_{w_{0}}^{-1},~~\lambda\in X_{*}.$
By [OpdamSpectral, Prop. 2.9], $^{\star}T_{w}=T_{w^{-1}}$ for $w\in W_{\mathrm{aff}}$. When $G$ is simply-connected, the involution $\dagger$ agrees with the Goldman involution of $\mathbf{H}$. Now we have
@@
  1. (a)
    We have $^{\dagger}C_{x}=(-1)^{\ell(\omega(x)_{f})}j(C_{w})=(-1)^{\ell(x)+\ell(\omega(x)_{f})}C^{\prime}_{x}$ for all $x\in W_{\mathrm{aff}}$.
  2. (b)
    We have $^{\dagger}\theta_{\lambda}=T_{w_{0}}\theta_{-\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}^{-1}T_{w_{0}}^{-1}$.
  3. (c)
    We have ${}^{\bullet}{}^{\star}{}^{\dagger}(-)={}^{*}(-)$ as involutions on $\mathbf{H}$.
  4. (d)
    We have $^{\bullet}(-)\circ{}^{\star}(-)=T_{w_{0}}^{-1}\bar{\kappa}(-)T_{w_{0}}$ as automorphisms of $\mathbf{H}$.
  5. (e)
    We have the commutative diagram
[…]
@@
By definition, if $x=\omega x^{\prime}$ for $x^{\prime}\in W\ltimes\mathbb{Z}\Phi^{\vee}$ and $\omega\in\pi_{1}(G)$, then $C^{\prime}_{x}=T_{\omega}C^{\prime}_{x^{\prime}}$ and $C_{x}=T_{\omega}C_{x^{\prime}}$. Further, if $\omega=\omega_{f}\omega_{t}\in W_{\mathrm{aff}}$ has $\ell(\omega)=0$, then $^{\dagger}T_{\omega}=(-1)^{\ell(\omega_{f})}T_{\omega}$. Therefore it suffices to show that
$\overline{{}^{\dagger}h}={}^{\dagger}\left(\bar{h}\right)=j(h)$
as $\mathcal{A}$-antilinear automorphisms of $\mathbf{H}$ for $G$ simply connected. In this case, $\ell(\lambda)\in 2\mathbb{Z}$ for any dominant $\lambda\in X_{*}$, whence $^{\dagger}(-)$ agrees with the involution $T_{w}\mapsto(-1)^{\ell(w)}\mathbf{q}^{\ell(w)}T_{w^{-1}}^{-1}$, $w\in W\ltimes\mathbb{Z}\Phi^{\vee}$. Therefore we compute
$\overline{\sum_{x}b_{x}{}^{\dagger}T_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\overline{\mathbf{q}^{\ell(x)}\bar{T}_{x}}=\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}T_{x}=j\left(\sum_{x}b_{x}T_{x}\right)$
@@
$^{\dagger}\left(\overline{\sum_{x}b_{x}T_{x}}\right)={}^{\dagger}\left(\sum_{x}\bar{b}_{x}(-1)^{\ell(x)}\mathbf{q}^{-\ell(x)}{}^{\dagger}T_{x}\right).$
Thus we have $\overline{{}^{\dagger}C_{w}}=(-1)^{\ell(w)}C^{\prime}_{w}$ for $G$ simply-connected, whence (a).
It suffices to prove (b) for $\lambda$ dominant, in which case
$^{\dagger}\theta_{\lambda}=\mathbf{q}^{-\frac{\ell(\lambda)}{2}}{}^{\dagger}T_{\lambda}=\mathbf{q}^{\frac{\ell(\lambda)}{2}}T_{-\lambda}^{-1}=\mathbf{q}^{\frac{\ell(\lambda)}{2}}\left({}^{\star}T_{\lambda}\right)^{-1}=\left({}^{\star}\theta_{\lambda}\right)^{-1}=\left(T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}\right)^{-1}=T_{w_{0}}\theta_{\bar{\kappa}(\lambda)}^{-1}T_{w_{0}}^{-1},$
where we used the equivalence of the definitions in Definition 6 (e). This shows (b).
For (c), we again compute, for $\lambda$ dominant, that
$^{\bullet}{}^{\star}{}^{\dagger}\theta_{\lambda}={}^{\bullet}{}^{\star}\left(T_{w_{0}}\theta_{-\bar{\kappa}(\lambda)}T_{w_{0}}^{-1}\right)={}^{\bullet}\theta_{-\lambda}=\theta_{-\lambda}.$
Agreement on $T_{w}$ for $w\in W$ follows from the fact that $^{\bullet\star}(-)$ is the identity on the finite Hecke algebra. This shows (c).
On the Bernstein subalgebra, part (d) follows from the definitions. On the finite Hecke algebra, we must show that the right hand side is the identity automorphism. For $s$ a finite reflection, write $w_{0}sw_{0}=\bar{\kappa}(s)$, again a finite reflection, so that $\bar{\kappa}(s)w_{0}s=w_{0}$, and
$T_{w_{0}}T_{s}T_{w_{0}}^{-1}=T_{w_{0}}T_{s}T_{s}^{-1}T_{w_{0}s}^{-1}=T_{w_{0}}T_{w_{0}s}^{-1}=T_{\bar{\kappa}(s)}T_{w_{0}s}T_{w_{0}s}^{-1}=T_{\bar{\kappa}(s)}.$
Therefore on the finite Hecke algebra, $\bar{\kappa}$ is given by conjugation by $T_{w_{0}}$.
Commutativity of the right square in (e) follows from Lemma 1, and commutativity of the left square follows from (c) and (d).
Given an $\mathbf{H}$ (or $\mathbf{H}_{\mathcal{K}}$)-module $M$, define $^{\dagger}M$ to be the same vector space with the $\mathbf{H}$-action twisted by this involution, and likewise for other automorphisms.
Corollary 1.
Twisting by $^{\dagger}(-)$ exchanges tempered and anti-tempered $H$-modules.
Proof.
Immediate from Lemma 2 (b). ∎
In fact, more is true: by [Kato, Thm. 2], $^{\dagger}M$ is the Aubert-Zelevinski dual of $M$.
Remark 4.
There is another natural auto-equivalence of the category of admissible representations of $G$ which exchanges tempered and anti-tempered representations, namely Bernstein’s cohomlogical duality. This functor differs from Aubert-Zelevinksi duality by the contragredient [SchneiderStuhler] (c.f. [PrasadNori]). On semisimple $H$-modules $\pi$, we have $^{\bar{\kappa}}\pi=\tilde{\pi}$, by [PrasadDuke, Prop. 6.3]. Therefore twist by the involution $^{*}(-)$, which manifestly exchanges tempered and anti-tempered representations, induces on semisimple modules the cohmological duality. However, it does not do so in general, because $\bar{\kappa}$ does not induce the contragredient in general. Indeed, for $G=\mathrm{SL}_{2}(F)$, $\bar{\kappa}=\mathrm{id}$ but not all non-unitary principal series of $G$ are self-dual. Therefore the operation $^{*}(-)$ is less natural than $^{\dagger}(-)$; this is perhaps reflected by the fact the formulas in Theorem 2 are nicer than those for $(\phi\circ{}^{*}(-))^{-1}(t_{w})$, which are related to those of the Theorem by Lemma 2 (e).
@@
There are bijections of sets
@@
where $K(u,s,\rho)$ is a standard module as in [KLDeligneLanglands]. Moreover, for a simple $J$-module $E$,
@@
    There is a unique two-sided cell $\mathbf{c}=\mathbf{c}(E)$ of $W_{\mathrm{aff}}$ such that $\mathrm{trace}\left({E}\,,{t_{w}}\right)\neq 0$ implies $w\in\mathbf{c}$.
@@
In particular, $\mathrm{trace}\left({E}\,,{t_{w}}\right)$ is independent of $\mathbf{q}$, and upon specializing $\mathbf{q}=q$ a prime power, will be a regular function in the twisting character in the setting of the Payley-Wiener theorem for the Iwahori-Hecke algebra of the $p$-adic group $G$, as we will explain in greater detail below.
We will comment even further in Section 2.4 on the necessity of twisting by some $\mathbf{H}$-involution exchanging tempered and anti-tempered modules, and how it is sufficient to twist $\phi$ by either $(-)^{*}$ or $(-)^{\dagger}$, but that the latter twist leads to nicer formulas.